Machinery is purchased on July 1 of the current fiscal year for $240,000. It is expected to have a useful life of four years, or 25,000 operating hours, and a residual value of $15,000. Compute the depreciation for the last six months of the current fiscal year ending December 31 by each of the following methods: a. Straight-line b. Double-declining-balance c. Units-of-activity (used for 1,600 hours during the current year)
Answer #1

Answers

Straight line= $28,150

DDB= $60,000

Units of activity = $14,400

Straight line Method

A

Cost

$ 2,40,000.00

B

Residual Value

$     15,000.00

C=A - B

Depreciable base

$ 2,25,000.00

D

Life [in years left ]

                         4

E=C/D

Annual SLM depreciation

$     56,250.00

Year

Book Value

Depreciation expense

Ending Book Value

Accumulated Depreciation

1

$ 2,40,000.00

$     28,125.00

$ 2,11,875.00

$ 28,125.00

.

.

Double declining Method

A

Cost

$ 2,40,000.00

B

Residual Value

$     15,000.00

C=A - B

Depreciable base

$ 2,25,000.00

D

Life [in years]

4

E=C/D

Annual SLM depreciation

$     56,250.00

F=E/C

SLM Rate

25.00%

G=F x 2

DDB Rate

50.00%

Year

Beginning Book Value

Depreciation rate

Depreciation expense

Ending Book Value

Accumulated Depreciation

1

$       2,40,000

50.00%

$           60,000

$   1,80,000

$ 60,000

.

.

Units of Production method

A

Cost

$ 2,40,000.00

B

Residual Value

$     15,000.00

C=A - B

Depreciable base

$ 2,25,000.00

D

Usage in units(in Miles)

25000

E

Depreciation per Mile

$                9.00

Year

Book Value

Usage

Depreciation expense

Ending Book Value

Accumulated Depreciation

1

$ 2,40,000.00

1600

$     14,400.00

$ 2,25,600.00

$ 14,400.00